OpenStudy (anonymous):

The perimeter of a rectangle is 46 inches. If the width of the rectangle is 11 inches, what is the length? 23 inches 12 inches 35 inches 24 inches

OpenStudy (anonymous):

i forgot what the formula was

OpenStudy (anonymous):

is it s + s + s idk

OpenStudy (anonymous):

it is a rectangle, it has two pairs of 4 sides. so, we can do 11 * 2 = 22 now, we have 2x + 22 = 46 46 - 22 = 24 24 / 2 = 12. the answer is 12.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem.